# Difficulty-Aware Agent Orchestration in LLM-Powered Workflows

### v2

@@ Introduction
Table 2: Performance comparison on the GAIA benchmark. Results are reported across three difficulty levels, with average scores shown in the last column. The best results are highlighted in bold.
@@
Table 3: Training, inference, and overall cost (in USD) on the MATH benchmark, along with corresponding accuracy. Our method achieves the lowest cost and highest accuracy. AFlow and MaAS use GPT-4o-mini, while other methods utilize an LLM pool.
Cost-effective.
@@
We emphasize the cost-efficiency of our agentic automation framework across two key dimensions: training expenditure and inference overhead. We compare against AFlow and MaAS, where AFlow represents the state-of-the-art (SOTA) among task-level frameworks, and MaAS is the SOTA among query-level frameworks. As shown in Table 3, AFlow incurs a substantial training cost of \$22.50 and inference cost of \$1.66, totaling \$24.16. In contrast, our method significantly reduces these costs to \$2.34 for training and \$0.27 for inference—only 10.4% and 16.3% of AFlow’s respective costs.
@@
This cost-efficiency is attributed to our adaptive model selection strategy, which dynamically leverages more affordable models such as LLaMA-3.1 or Qwen-2-72B when sufficient, rather than defaulting to high-cost models like GPT-4o-mini. Notably, our method not only reduces cost but also achieves the highest accuracy of 55.37%, outperforming both AFlow and MaAS.
MasRouter adopts a collaborative paradigm, assigning multiple LLMs to role-play in handling a query. However, it suffers from two drawbacks: (1) redundant participation of LLMs in each collaborative step, and (2) lack of adaptation for easy queries, leading to excessive cost without proportional performance gains.
Case Study
Figure 2 presents visual comparisons of workflows produced by our framework and MaAS on MATH, HumanEval, and GAIA. For easy queries (middle), our method simplifies the workflow without compromising accuracy, thereby reducing cost. In contrast, for difficult queries (left), MaAS generates oversimplified workflows that fail to solve the problem, while our framework assembles a more complex, heterogeneous workflow that effectively combines different models’ strengths—achieving higher performance with controlled cost (\$0.0035 vs. \$0.0028 for MaAS).
Furthermore, under the same workflow structure, our method achieves comparable accuracy to MaAS while incurring lower cost, highlighting the effectiveness of our query-aware complexity adaptation and fine-grained cost–performance control.
Figure 3 shows the distribution of LLMs selected by DAAO on the MMLU and MATH benchmarks. Our framework selects LLMs based on both domain and difficulty, effectively exploiting the specialized capabilities of each model.
@@
Figure 3: Distribution of LLM selections made by DAAO on the MATH and MMLU benchmarks. The results reflect how the framework dynamically routes queries to different LLMs based on domain and difficulty.
Dataset HumanEval MATH
@@
Vanilla $93.37$ $1.23$ $55.37$ $0.55$
w/o DA $90.21$ $1.34$ $50.18$ $0.67$
w/o LS $92.69$ $1.60$ $53.42$ $0.89$
w/o $\textbf{C}(\cdot)$ $93.21$ $1.58$ $55.40$ $1.00$
Table 4: Ablation study of DAAO on HumanEval and MATH. We report performance (Pass@1 or Accuracy) and corresponding inference cost. w/o DA removes the difficulty-aware module; w/o LS disables LLM selection; w/o $\textbf{C}(\cdot)$ omits the cost-awareness component.
Ablation Study
We conduct an ablation study on two key components of our DAAO framework: (1) w/o DA, removing the difficulty-aware module; and (2) w/o LS, removing the LLM selector and routing all subtasks to a fixed LLM;(3) w/o $\textbf{C}(\cdot)$, eliminating the cost constraint in Equation 5. As shown in Table 4, removing the difficulty-aware module leads to the largest drop in both accuracy and efficiency, especially on the MATH dataset. This highlights the importance of adaptive reasoning control based on estimated query difficulty (e.g., dynamically adjusting the number of reasoning layers instead of using a fixed number).
Removing the LLM router slightly affects accuracy, but leads to a notable increase in inference cost, as it prevents the system from using lightweight models when appropriate. Removing $\textbf{C}(\cdot)$ does not significantly impact the performance, but it disrupts the adaptive capability of DAAO to query difficulty. Overall, the results demonstrate that both components are crucial for balancing performance and cost in multi-step reasoning tasks.
Conclusion

### v3

@@ 4.2. Performance Analysis
Table 2. Performance comparison on the GAIA benchmark. Results are reported across three difficulty levels, with average scores shown in the last column. The best results are highlighted in bold.
@@
Table 3. Training, inference, and overall cost (in USD) on the MATH benchmark, along with corresponding accuracy. Our method achieves the lowest cost and highest accuracy. AFlow and MaAS use GPT-4o-mini, while other methods utilize an LLM pool.
@@
4.2.2. Cost-effective.
@@
We emphasize the cost-efficiency of our agentic automation framework across two key dimensions: training expenditure and inference overhead. We compare against AFlow and MaAS, where AFlow represents the state-of-the-art (SOTA) among task-level frameworks, and MaAS is the SOTA among query-level frameworks. As shown in Table 3, AFlow incurs a substantial training cost of \$22.50 and inference cost of \$1.66, totaling \$24.16. In contrast, our method significantly reduces these costs to \$2.34 for training and \$0.27 for inference—only 10.4% and 16.3% of AFlow’s respective costs. MasRouter adopts a collaborative paradigm, assigning multiple LLMs to role-play in handling a query. However, it suffers from two drawbacks: (1) redundant participation of LLMs in each collaborative step, and (2) lack of adaptation for easy queries, leading to excessive cost without proportional performance gains. Notably, our method not only reduces cost but also achieves the highest accuracy of 55.37%, outperforming both AFlow and MaAS.
@@
This cost-efficiency is attributed to two strategies: (1) Our difficulty-awareness strategy, which generates adaptive workflows for queries, employing simple workflows for easy queries while allocating more resources to the generation of workflows for more complex queries; (2) Our adaptive model selection strategy, which dynamically leverages more affordable models (such as LLaMA-3.1 or Qwen-2-72B) when sufficient, rather than defaulting to high-cost models like GPT-4o-mini.
Figure 2. The visualization of the workflow generated by DAAO. Colors represent different models assigned to each operator.
4.3. Case Study
As shown in Figure 2, DAAO generates different workflows for queries of varying difficulty. For simple queries, it produces streamlined workflows, sometimes using only a single operator. For medium or difficult queries, it constructs deeper and more complex workflows by exploring a broader combination of operators. This demonstrates DAAO’s difficulty-aware paradigm: selecting economical workflows for simple queries to enable rapid completion, while leveraging sophisticated workflows for complex queries to meet higher demands.
Table 4. Cross-domain optimization performance
Train on Test on Perf.
MATH MATH 55.37
MATH GSM8K 95.44
MATH+GSM8K MATH 56.42
MATH+GSM8K GSM8K 95.70
HumanEval HumanEval 94.65
HumanEval MATH 54.46
HumanEval+MATH HumanEval 95.00
HumanEval+MATH MATH 55.50
$4.$ Inductive Ability Analysis
4.4.1. Cross-domain Optimization.
We present the performance results of our cross-domain training experiments (see Table 4), evaluating the impact of multi-domain joint optimization on generalization. We observe that single-domain training achieves baseline performance on the MATH dataset, with strong transfer to GSM8K, primarily due to their shared mathematical reasoning requirements. Moreover, since MATH problems are generally more challenging than those in GSM8K, training on MATH equips the framework with advanced problem-solving skills that effectively generalize to the relatively simpler GSM8K. In contrast, we find that single-domain training on HumanEval yields high fidelity on code generation (94.65%) but limited transfer to MATH (54.46%), underscoring domain-specific overfitting. Notably, we achieve modest improvements across target tasks (approximately 0.35%–1.05%) through joint training (e.g., MATH+GSM8K and HumanEval+MATH), without inducing catastrophic forgetting, indicating that our simultaneous exposure to multiple domains promotes shared representation learning and enhances model robustness.
Overall, our multi-domain setup not only preserves intra-domain proficiency but also slightly boosts cross-domain generalization. Furthermore, within the same domain, training on more complex benchmark datasets can further enhance the framework’s ability to generate workflows for complex queries.
@@ 4.3. Case Study
Figure 3. Distribution of LLM selections made by DAAO on the MATH and MMLU benchmarks.
4.4.2. LLM Router Analysis.
In this section, we validate that DAAO does not exhibit a preference for any particular LLM and demonstrate its ability to generalize well to unseen LLMs without requiring extensive pretraining. Figure 3 illustrates the distribution of LLMs selected by DAAO on the MATH and MMLU datasets before and after the addition of DeepSeek-v3, with the new model being chosen 29% and 15% of the time, respectively. By intelligently selecting models that match the difficulty and domain of each query, DAAO improved the accuracy on MATH from 55.37% to 56.20% and increased the accuracy on MMLU from 84.90% to 85.66%.
Table 5. Ablation study of DAAO on HumanEval and MATH. We report performance (Pass@1 or Accuracy) and corresponding inference cost. w/o DA removes the difficulty-aware module; w/o LS disables LLM selection; w/o $\textbf{C}(\cdot)$ omits the cost-awareness component.
Dataset
HumanEval MATH
@@
Vanilla
$94.65$ $1.10$ $55.37$ $0.55$
w/o DA $92.21$ $1.64$ $52.18$ $0.88$
w/o LS $92.69$ $1.38$ $53.24$ $0.79$
w/o $\textbf{C}(\cdot)$ $94.72$ $1.88$ $55.40$ $1.00$
4.5. Framework Analysis
4.5.1. Ablation Study.
We conduct an ablation study on three key components of our DAAO framework: (1) w/o DA, removing the difficulty-aware module; and (2) w/o LS, removing the LLM selector and routing all subtasks to a fixed LLM;(3) w/o $\textbf{C}(\cdot)$, eliminating the cost constraint in Equation 5. As shown in Table 5, removing the difficulty-aware module leads to the largest drop in both accuracy and efficiency, especially on the MATH dataset. This highlights the importance of adaptive reasoning control based on estimated query difficulty (e.g., dynamically adjusting the number of reasoning layers instead of using a fixed number). Removing the LLM router slightly affects accuracy, but leads to a notable increase in inference cost, as it prevents the system from using lightweight models when appropriate. Removing $\textbf{C}(\cdot)$ does not significantly impact the performance, but it disrupts the adaptive capability of DAAO to query difficulty. Overall, the results demonstrate that both components are crucial for balancing performance and cost in multi-step reasoning tasks.
Figure 4. Sensitivity analysis of DAAO on HumanEval. The unit of cost per query (right) and performance (left) is $10^{-3}$ · \$ and pass@1 (%), respectively.
Table 6. Performance(%) and Average Cost($10^{-3}\mathdollar$) across thresholds $P$ on HumanEval and GSM8K.
Dataset Metric 0.1 0.2 0.3 0.4 0.5 0.6 0.7
HumanEval Perf. 92.80 93.60 94.65 94.42 94.80 94.77 94.80
Cost 0.86 0.94 1.10 1.29 1.28 1.78 2.30
GSM8K Perf. 91.99 93.46 94.40 94.34 94.70 94.75 94.40
Cost 0.40 0.45 0.50 0.59 0.68 0.77 0.96
4.5.2. Sensitivity Analysis.
We analyzed the sensitivity of DAAO to four core parameters: the maximum number of layers in the agentic supernet $L_{max}$ in Equation Equation 14, the cost penalty coefficient $\lambda$ in Equation 5, the sampling count $K$ and the threshold $\tau$ in Equation Equation 17. The results are shown in Figure 4 and Table 6. For the parameter $L_{max}$, we observed a significant performance improvement when $L_{max}$ increased from 4 to 5 (from 92.9% to 94.6%). However, further increases in $L_{max}$ only yielded marginal performance gains while significantly increasing the inference cost per query. Considering both performance and cost, we selected $L_{max}$ = 5. For the parameter $\lambda$, we found that larger $\lambda$ values led DAAO to favor more cost-efficient solutions, but with a slight performance degradation. For the parameter $K$, we note that performance is suboptimal with highest variance when $K$ = 2. Increasing $K$ to 4 effectively achieves a satisfactory low-variance estimation. For the threshold $\tau$, as shown in Table 6, performance improves as the threshold $\tau$ increases. However, the gain stops growing beyond 0.3. At the same time, a higher threshold $\tau$ raises inference costs because more operators are activated in each layer. Therefore, we set the threshold $\tau$ to 0.3 to balance performance and efficiency.
5. Conclusion
@@
Appendix A Notations
  • O=M,SO={M,S}: An agentic operator defined by pairing an LLM $M$ with a protocol $S$.
  • $M$: A large language model (LLM) instance.
  • $\mathbb{M}$: The set of all available LLMs.
  • $S$: A collaboration or reasoning protocol (e.g., Chain-of-Thought, Debate).
  • $\mathbb{S}$: The set of all feasible protocols.
  • $\mathbb{O}\subseteq\mathbb{M}\times\mathbb{S}$: The catalog of feasible agentic operators.
  • $G={\mathcal{V},\mathcal{E}}$: A directed acyclic graph (DAG) representing an agentic workflow, with operator nodes $\mathcal{V}$ and dependency edges $\mathcal{E}$.
  • $\mathcal{V}=\bigcup_{\ell=1}^{L}\mathcal{V}{\ell}$: The layered decomposition of workflow nodes across $L$ layers.
  • $L$: The number of workflow layers (depth).
  • $\mathcal{A}\;=\;\big\{\{\pi_{l}(O)\}_{O\in\mathbb{O}}\big\}_{l=1}^{L}$: Layered operator-selection policy, where $\pi_{\ell}(O)$ is the probability of selecting operator $O$ at layer $\ell$.
  • $\pi^{(L)}(L\mid\mathcal{Q},z)$: Distribution over workflow depth given $\mathcal{Q}$ and $z$.
  • $\pi^{(O)}_{l}(O\mid\mathcal{Q},l,z)$: Distribution over operators at layer $\ell$ conditioned on $\mathcal{Q}$ and difficulty embedding $z$.
  • $\pi^{(M)}(M\mid\mathcal{Q},O,z)$: Model-routing policy assigning an LLM $M$ to operator $O$ based on $\mathcal{Q}$ and $z$.
  • $z\in\mathbb{R}^{k}$: Latent difficulty embedding of the query produced by the difficulty estimator $N{\theta_{d}}$.
  • $d\in(0,1)$: Scalar difficulty score decoded from $z$, where larger $d$ indicates a harder query.
  • $N_{\theta_{d}}$: Difficulty estimator (a VAE with a learned difficulty head).
  • $N_{\theta_{o}}$: Operator allocator that selects operators per layer.
  • $N_{\theta_{m}}$: Cost-aware LLM router that selects models for operators.
  • $N_{\theta_{L}}$: Workflow depth selector determining $L$ based on $z$.
  • $U(G;\mathcal{Q},a)$: Utility (e.g., accuracy) achieved by executing workflow $G$ on query $\mathcal{Q}$.
  • $C(G;\mathcal{Q})$: Inference cost (e.g., token usage, latency) of workflow $G$ for query $\mathcal{Q}$.
  • $\lambda$: Trade-off coefficient balancing utility and cost in Eq. (5).
  • $v(\cdot)$: Text or operator embedding function (e.g., MiniLM, SBERT).
  • $\tau$: Cumulative evidence threshold controlling layer width (Eq. (17)).
  • $K$: The number of samples in each data set.
Appendix B Technical Details
B.1. Operator Space
In this section, we detail the initialization of operator nodes as follows:
  1. (1)
    Chain-of-Thought (CoT). CoT (Wei et al., 2022) reasoning encourages the LLM to think step by step rather than directly outputting an answer. This approach enhances its capability to solve complex problems through intermediate reasoning steps, improving task handling and providing greater transparency in the decision-making process.
  2. (2)
    LLM-Debate. LLM-Debate (Du et al., 2023) allows multiple LLMs to debate, leveraging diverse perspectives to identify better solutions. In practice, we initialize three debaters and permit up to two debate rounds.
  3. (3)
    Self-Consistency. Adopting the methodology from (Wang et al., 2023b), this operator aggregates five CoT reasoning paths and determines the final answer through majority voting.
  4. (4)
    Self-Refine. Following (Madaan et al., 2023), this operator initially generates an answer using CoT reasoning, then prompts the agent to self-reflect iteratively. We set a maximum of five refinement iterations.
  5. (5)
    Ensemble. Inspired by LLM-Blender (Jiang et al., 2023), this operator involves three LLM-powered agents from different sources outputting answers to the same query. The pairwise ranking is used to evaluate and aggregate their responses into a final solution.
  6. (6)
    Testing. Following the test designer in AgentCoder (Huang et al., 2023), this operator is used for generating test cases for the generated code.
  7. (7)
    ReAct. Following (Yao et al., 2023), this operator enables the agent to leverage versatile tools, including code interpreter, web searching, external knowledge database, etc., to handle diverse user demands.
B.2. Embedding Function
Following established practices (Feng et al., 2024), we first employ an LLM to generate a comprehensive profile description for each operator. Subsequently, a lightweight text embedding model (in our case, MiniLM (Wang et al., 2020)) is used to encode the profile into a fixed-dimensional embedding. The prompt for generating the operator profile is as follows:
Embedding Prompt prompt = """You are a highly proficient expert in designing and defining operators for large language models (LLMs). Your primary objective is to meticulously generate the description and interface fields for a specified operator based on its provided Python implementation. The generated content must be accurate, efficient, and precisely reflect the functionality of the operators code. To ensure consistency, quality, and adherence to best practices, refer to the following examples of previously defined operators: { "Generate": { "description": "Generates anything based on customized input and instruction.", "interface": "generate(input: str, instruction: str) -> dict with key response of type str" }, "ScEnsemble": { "description": "Uses self-consistency to select the solution that appears most frequently in the solution list, improving the selection to enhance the choice of the best solution.", "interface": "sc_ensemble(solutions: List[str], problem: str) -> dict with key response of type str" } } Now, given the following operator code. This code encompasses the function signature, parameters with type annotations, internal logic, and return statements essential for comprehensively understanding the operators purpose and behavior.Please provide its description and interface fields in the same format. [operator code] """
Appendix C Experimental Details
In this section, we introduce each dataset along with its primary evaluation metric.
  • HumanEval: uses Pass@k as the primary metric to measure the proportion of correctly generated samples in code generation tasks.
  • MBPP: also uses Pass@k to evaluate the model’s ability to generate solutions for programming problems.
  • GSM8K: uses Accuracy to assess the model’s correctness in mathematical reasoning problems.
  • MATH: uses Accuracy as the core metric to evaluate the model’s performance on solving mathematics problems.
  • MMLU: uses Accuracy to measure the model’s performance on multi-domain knowledge question-answering tasks.
  • GAIA: uses Accuracy as the primary metric to evaluate the model’s performance on general AI tasks.